OpenStudy (anonymous):

find the distance between the two points. round to the nearest tenth if neccessary.(2,5) and (-1,-5)

OpenStudy (btaylor):

\[d = \sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}\]So, by the distance formula,\[d = \sqrt{(-1 - 2)^2 + (-5 - 5)^2} = \sqrt{(-3)^2+(-10)^2} = \sqrt{109} \approx 10.4\]

OpenStudy (anonymous):

ok how di u get 109

OpenStudy (btaylor):

\[(-3)^2 + (-10)^2 = 9 + 100 = 109\]

OpenStudy (anonymous):

ok thanks but i dont think i under stand how u got 9 either

OpenStudy (btaylor):

\[(-3)^2 = -3 \times -3 = 9\]